# How to Scale Inverse RL to Large State Spaces? A Provably Efficient Approach

### v1

@@ 1 Introduction
In practical applications, the chosen IRL formulation has to be tackled by algorithms that use a finite number of demonstrations and a limited knowledge of the dynamics of the environment. In the common online IRL scenario, the learner explores the (unknown) environment, and exploits this additional information to improve its performance on the IRL task [e.g., 37, 31, 36, 63, 29]. On this basis, the IRL approach ($iii$) based on the feasible set [37, 36] displays desirable properties since “postpones” the choice of the heuristic and/or enforcement of additional constraints, with the advantage of analyzing the intrinsic complexity of the IRL problem only, without being obfuscated by other factors. In other words, this recent formulation of the IRL problem paves the way for the design and analysis of provably efficient IRL algorithms, endowed with solid theoretical guarantees.
However, the algorithms designed for learning the feasible set currently available in the literature [e.g., 37, 31, 36, 63, 29] struggle when attempting to scale them to IRL problems with large state spaces. This is apparent because their sample complexity exhibits an explicit dependence on the cardinality of the state space. This inevitably represents a major limitation since most real-world scenarios concern problems with large, or even continuous, state spaces [e.g., 14, 7, 38, 13].
In this context, function approximation represents an essential tool to tackle the curse of dimensionality and enforce generalization [50, 39]. Linear Markov Decision Processes (MDPs) [22, 62] offer a simple but powerful structure, in which we assume the reward function and the transition model can be expressed as linear combinations of known features, that permits theoretical analysis of the sample complexity. Even though many extensions have been developed [59, 21, 12], the Linear MDPs framework typically represents one of the first function approximation settings to analyze when focusing on a novel problem, before moving to more complex settings [e.g., 58, 56].
@@ 3 Limitations of the Feasible Set
The Feasible Set.  According to the standard definition [e.g., 37, 31, 36, 63, 29], the feasible set contains the rewards that make the expert’s policy $\pi^{E}$ optimal, as defined below.
@@
Without function approximation, the feasible set contains a variety of rewards for any deterministic policy. In Linear MDPs, due to the feature map, the feasible set might exhibit some degeneracy. [^3] We exemplify this proposition in Appendix B.1. In Appendix B.4 we generalize to infinite state spaces.
@@
where $\mathcal{A}^{E}_{h}(s)\coloneqq\{a\in\mathcal{A}|\pi^{E}_{h}(\cdot|s)>0\}$ for every $s\in\mathcal{S}$. If for none of the $H$ pairs of sets $(\Phi_{h}^{\pi^{E}},\overline{\Phi}_{h})$ there exists a separating hyperplane, then $\mathcal{R}_{p,\pi^{E}}=\{0\}$, i.e., the feasible set with linear rewards in $\phi$ contains only the reward function that assigns zero reward to all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\llbracket H\rrbracket$.
Intuitively, expert’s actions must have the largest optimal $Q$-value among all actions, and linearity imposes the “separability” requirement. The result holds also for MDPs with linear rewards only.
@@ 4.3 A Learning Framework for Online IRL Classification
Remark 4.2.
For $\eta\geq 0$, let $\mathcal{R}_{\eta}\coloneqq\{r\in\mathcal{R}\,|\,\overline{\mathcal{C}}_{p,\pi^{E}}(r)\leq\eta\}$ and ^$\widehat{\mathcal{R}}_{\eta}\coloneqq\{r\in\mathcal{R}\,|\,\widehat{\mathcal{C}}(r)\leq\eta\}$ denote the sets of rewards positively classified using, respectively, the true (non)compatibility $\overline{\mathcal{C}}_{p,\pi^{E}}$ and the estimate $\widehat{\mathcal{C}}$ constructed by an $(\epsilon,\delta)$-PAC algorithm. Then, with probability $1-\delta$, it holds that: ^$\widehat{\mathcal{R}}_{\Delta-\epsilon}\subseteq\mathcal{R}_{\Delta}\subseteq\widehat{\mathcal{R}}_{\Delta+\epsilon}$. Thus, we can trade-off the amount of “false negatives” (resp. “false positives”) by, e.g., choosing the threshold $\Delta\leftarrow\Delta+\epsilon$ (resp. $\Delta\leftarrow\Delta-\epsilon$).
Figure 1: Flow-chart of CATY-IRL.
@@ 5 CATY-IRL: A Provably Efficient Algorithm for IRL
Classification phase.  The classification performs the estimation $\widehat{\mathcal{C}}(r)$ of the (non)compatibility term $\overline{\mathcal{C}}_{p,\pi^{E}}(r)$ for the single input reward $r\in\mathcal{R}$ by splitting it into two independent estimates: $\widehat{J}^{E}(r)\approx J^{\pi^{E}}(r;p)$, which is computed with $\mathcal{D}^{E}$ only, and $\widehat{J}^{*}(r)\approx J^{*}(r;p)$, which is computed with $\mathcal{D}$ only. Concerning $\widehat{J}^{E}(r)$, when the reward is linear $r_{h}(s,a)=\langle\phi(s,a),\theta_{h}\rangle$, CATY-IRL uses $\mathcal{D}^{E}$ to construct an empirical estimate $\widehat{\psi}^{E}\approx\psi^{p,\pi^{E}}$ of the expert’s expected feature count [6]. Otherwise, it directly estimates $\widehat{d}^{E}\approx d^{p,\pi^{E}}$ the expert’s occupancy measure. Such estimates can be used to derive $\widehat{J}^{E}(r)$ straightforwardly. Regarding $\widehat{J}^{*}(r)$, CATY-IRL exploits the planning phase of the corresponding RFE (or BPI) algorithm adopted at exploration phase. [^8] RFE/BPI algorithms, at planning phase, return a policy, and not its estimated performance. Since BPI-UCBVI, RF-Express, and RFLin each compute an estimate of $J^{*}(r;p)$ as an intermediate step, with negligible abuse of notation, we assume that they output such estimate. Finally, CATY-IRL applies the (potentially negative) input threshold $\Delta$ to the difference $\widehat{J}^{*}(r)-\widehat{J}^{E}(r)$ to perform the classification. See Appendix D for the full pseudo-code. Clearly, CATY-IRL can be implemented in practice, since it considers a single reward at a time instead of computing the full feasible set.
@@ Appendix B Additional Results and Proofs for Section 3
Let $\mathcal{M}=(\mathcal{S},\mathcal{A},H,d_{0})\cup\{\pi^{E}\}$ be an IRL instance such that $\mathcal{S}=\{s_{1},s_{2}\},\mathcal{A}=\{a_{1},a_{2}\},H=1,d_{0}(s_{1})=d_{0}(s_{2})=1/2,\pi^{E}(s_{1})=\pi^{E}(s_{2})=a_{1}$. Consider the feature mapping $\phi_{1}$ s.t. $\phi_{1}(s,a)=\mathds{1}\{a=a_{1}\}$ for all $s\in\mathcal{S}$. Then, we have $\Phi^{\pi^{E}}=\{1\}$ and $\overline{\Phi}=\{0\}$. Clearly, these sets can be separated by any hyperplane $w\in\mathbb{R}_{>0}$, since $1\cdot w>0\cdot w$, and so $\mathcal{R}_{p,\pi^{E}}\neq\{0\}$ (actually, $\mathcal{R}_{p,\pi^{E}}=(0,1]$).
@@
Consider the same IRL instance as in the previous example, but this time consider the feature mapping $\phi_{2}$ s.t. $\phi_{2}(s_{1},a)=\mathds{1}\{a=a_{1}\}$, and $\phi_{2}(s_{2},a)=\mathds{1}\{a=a_{2}\}$. Then, we have $\Phi^{\pi^{E}}=\{0,1\}$ and $\overline{\Phi}=\{0,1\}$. Clearly, the two sets coincide, thus they cannot be separated, and $\mathcal{R}_{p,\pi^{E}}=\{0\}$.

### v2

# How does Inverse RL Scale to Large State Spaces? A Provably Efficient Approach

@@ 1 Introduction
In practical applications, the chosen IRL formulation has to be tackled by algorithms that use a finite number of demonstrations and a limited knowledge of the dynamics of the environment. In the common online IRL scenario, the learner explores the (unknown) environment, and exploits this additional information to improve its performance on the IRL task [e.g., 39, 33, 38, 68, 31]. On this basis, the IRL approach ($iii$) based on the feasible set [39, 38] displays desirable properties since “postpones” the choice of the heuristic and/or enforcement of additional constraints, with the advantage of analyzing the intrinsic complexity of the IRL problem only, without being obfuscated by other factors. In other words, this recent formulation of the IRL problem paves the way for the design and analysis of provably efficient IRL algorithms, endowed with solid theoretical guarantees.
However, the algorithms designed for learning the feasible set currently available in the literature [e.g., 39, 33, 38, 68, 31] struggle when attempting to scale them to IRL problems with large state spaces. This is apparent because their sample complexity exhibits an explicit dependence on the cardinality of the state space. This inevitably represents a major limitation since most real-world scenarios concern problems with large, or even continuous, state spaces [e.g., 15, 7, 40, 14].
In this context, function approximation represents an essential tool to tackle the curse of dimensionality and enforce generalization [54, 41]. Linear Markov Decision Processes (MDPs) [23, 67] offer a simple but powerful structure, in which we assume the reward function and the transition model can be expressed as linear combinations of known features, that permits theoretical analysis of the sample complexity. Even though many extensions have been developed [64, 22, 13], the Linear MDPs framework typically represents one of the first function approximation settings to analyze when focusing on a novel problem, before moving to more complex settings [e.g., 63, 61].
@@ 3 Limitations of the Feasible Set
The Feasible Set.  According to the standard definition [e.g., 39, 33, 38, 68, 31], the feasible set contains the rewards that make the expert’s policy $\pi^{E}$ optimal, as defined below.
@@
Without function approximation, the feasible set contains a variety of rewards for any deterministic policy. In Linear MDPs, due to the feature map, the feasible set might exhibit some degeneracy. [^3] We exemplify this proposition in Appendix B.1. In Appendix B.4 we generalize to infinite state spaces. Definition 3.1 can be adapted to Linear MDPs with feature map $\phi$ as: $\mathcal{R}_{\phi,p,\pi^{E}}\coloneqq\{r\in\mathfrak{R}\,|\,J^{\pi^{E}}(r;p)=J^{*}(r;p)\wedge\exists\theta:\llbracket H\rrbracket\to\mathbb{R}^{d},\forall(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\llbracket H\rrbracket:\,r_{h}(s,a)=\langle\phi(s,a),\theta_{h}\rangle\}$. We omit $\phi$ in $\mathcal{R}_{\phi,p,\pi^{E}}$ for notational simplicity.
@@
where $\mathcal{A}^{E}_{h}(s)\coloneqq\{a\in\mathcal{A}|\pi^{E}_{h}(\cdot|s)>0\}$ for every $s\in\mathcal{S}$. If for none of the $H$ pairs of sets $(\Phi_{h}^{\pi^{E}},\overline{\Phi}_{h})$ there exists a separating hyperplane, then $\mathcal{R}_{p,\pi^{E}}=\{\overline{r}\}$, with $\overline{r}_{h}(s,a)=0$ $\forall(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\llbracket H\rrbracket$ i.e., the feasible set with linear rewards in $\phi$ contains only the reward function that assigns zero reward everywhere.
Intuitively, expert’s actions must have the largest optimal $Q$-value among all actions, and linearity imposes the “separability” requirement. The result holds also for MDPs with linear rewards only. We exemplify Proposition 3.1 in Appendix B.1.
@@ 4.1 Compatible Rewards
For a discussion on comparing the (non)compatibility of different rewards, see Appendix C.4.
@@ 4.3 A Learning Framework for Online IRL Classification
Remark 4.2.   For $\eta\geq 0$, let $\mathcal{R}_{\eta}\coloneqq\{r\in\mathcal{R}\,|\,\overline{\mathcal{C}}_{p,\pi^{E}}(r)\leq\eta\}$ and ^$\widehat{\mathcal{R}}_{\eta}\coloneqq\{r\in\mathcal{R}\,|\,\widehat{\mathcal{C}}(r)\leq\eta\}$ denote the sets of rewards positively classified using, respectively, the true (non)compatibility $\overline{\mathcal{C}}_{p,\pi^{E}}$ and the estimate $\widehat{\mathcal{C}}$ constructed by an $(\epsilon,\delta)$-PAC algorithm. Then, with probability $1-\delta$, it holds that: ^$\widehat{\mathcal{R}}_{\Delta-\epsilon}\subseteq\mathcal{R}_{\Delta}\subseteq\widehat{\mathcal{R}}_{\Delta+\epsilon}$. Thus, we can trade-off the amount of “false negatives” (resp. “false positives”) by, e.g., choosing the threshold $\Delta\leftarrow\Delta+\epsilon$ (resp. $\Delta\leftarrow\Delta-\epsilon$).
@@ 5 CATY-IRL: A Provably Efficient Algorithm for IRL
Classification phase.  The classification performs the estimation $\widehat{\mathcal{C}}(r)$ of the (non)compatibility term $\overline{\mathcal{C}}_{p,\pi^{E}}(r)$ for the single input reward $r\in\mathcal{R}$ by splitting it into two independent estimates: $\widehat{J}^{E}(r)\approx J^{\pi^{E}}(r;p)$, which is computed with $\mathcal{D}^{E}$ only, and $\widehat{J}^{*}(r)\approx J^{*}(r;p)$, which is computed with $\mathcal{D}$ only. Concerning $\widehat{J}^{E}(r)$, when the reward is linear $r_{h}(s,a)=\langle\phi(s,a),\theta_{h}\rangle$, CATY-IRL uses $\mathcal{D}^{E}$ to construct an empirical estimate $\widehat{\psi}^{E}\approx\psi^{p,\pi^{E}}$ of the expert’s expected feature count [6]. Otherwise, it directly estimates $\widehat{d}^{E}\approx d^{p,\pi^{E}}$ the expert’s occupancy measure. Such estimates can be used to derive $\widehat{J}^{E}(r)$ straightforwardly. Regarding $\widehat{J}^{*}(r)$, CATY-IRL exploits the planning phase of the corresponding RFE (or BPI) algorithm adopted at exploration phase. [^8] RFE/BPI algorithms, at planning phase, return a policy, and not its estimated performance. Since BPI-UCBVI, RF-Express, and RFLin each compute an estimate of $J^{*}(r;p)$ as an intermediate step, with negligible abuse of notation, we assume that they output such estimate. Finally, CATY-IRL applies the (potentially negative) input threshold $\Delta$ to the difference $\widehat{J}^{*}(r)-\widehat{J}^{E}(r)$ to perform the classification. See Appendix D for the full pseudo-code. Clearly, CATY-IRL can be implemented in practice, since it considers a single reward at a time instead of computing the full feasible set, and it is computationally efficient in linear MDPs, since it uses a computationally efficient algorithm as subroutine (see [62]).
@@ 7 Conclusions
Acknowledgments and Disclosure of Funding
AI4REALNET has received funding from European Union’s Horizon Europe Research and Innovation programme under the Grant Agreement No 101119527. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
Funded by the European Union - Next Generation EU within the project NRPP M4C2, Investment 1.,3 DD. 341 - 15 march 2022 - FAIR - Future Artificial Intelligence Research - Spoke 4 - PE00000013 - D53C22002380006.
@@ Appendix B Additional Results and Proofs for Section 3
Let $\mathcal{M}=(\mathcal{S},\mathcal{A},H,d_{0})\cup\{\pi^{E}\}$ be an IRL instance such that $\mathcal{S}=\{s_{1},s_{2}\},\mathcal{A}=\{a_{1},a_{2}\},H=1,d_{0}(s_{1})=d_{0}(s_{2})=1/2,\pi^{E}(s_{1})=\pi^{E}(s_{2})=a_{1}$. Consider the feature mapping $\phi_{1}$ s.t. $\phi_{1}(s,a)=\mathds{1}\{a=a_{1}\}$ for all $s\in\mathcal{S}$. Then, we have $\Phi^{\pi^{E}}=\{1\}$ and $\overline{\Phi}=\{0\}$. Clearly, these sets can be separated by any hyperplane $w\in\mathbb{R}_{>0}$, since $1\cdot w>0\cdot w$, and so $\mathcal{R}_{p,\pi^{E}}\neq\{\overline{r}\}$, with $\overline{r}_{h}(s,a)=0$ $\forall(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\llbracket H\rrbracket$. Actually, $\mathcal{R}_{p,\pi^{E}}=\{r\in\mathfrak{R}\,|\,\exists\theta\in(0,1]:\,r_{1}(s,a)=\langle\phi(s,a),\theta\rangle\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}\}$.
@@
Consider the same IRL instance as in the previous example, but this time consider the feature mapping $\phi_{2}$ s.t. $\phi_{2}(s_{1},a)=\mathds{1}\{a=a_{1}\}$, and $\phi_{2}(s_{2},a)=\mathds{1}\{a=a_{2}\}$. Then, we have $\Phi^{\pi^{E}}=\{0,1\}$ and $\overline{\Phi}=\{0,1\}$. Clearly, the two sets coincide, thus they cannot be separated, and $\mathcal{R}_{p,\pi^{E}}=\{\overline{r}\}$, with $\overline{r}_{h}(s,a)=0$ $\forall(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\llbracket H\rrbracket$.
@@ Conclusions.
C.4 Comparing the (non)compatibility of various rewards
In Section 4, we said that rewards $r$ with smaller values of $\overline{\mathcal{C}}_{p,\pi^{E}}(r)$ are more compatible with $\pi^{E}$ in $\mathcal{M}=(\mathcal{S},\mathcal{A},H,d_{0},p)$. However, one might provide the following “counter-example”:
Example C.1 (Question by Reviewer KyLX).
Let $r^{1},r^{2}$ be two rewards such that $r^{2}_{h}(s,a)=2r^{1}_{h}(s,a)\geq 0$ for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\llbracket H\rrbracket$. Then, clearly, $\overline{\mathcal{C}}_{p,\pi^{E}}(r^{2})=2\overline{\mathcal{C}}_{p,\pi^{E}}(r^{1})$. Therefore, based on Section 4, we say that reward $r^{1}$ is more compatible than $r^{2}$ w.r.t. $\pi^{E}$ in $\mathcal{M}$. However, since $r^{2}$ is just $r^{1}$ re-scaled by a constant, the two MDPs $\mathcal{M}\cup\{r^{1}\}$ and $\mathcal{M}\cup\{r^{2}\}$ should be “equivalent”, thus, $r^{1}$ and $r^{2}$ should be, intuitively, equally compatible with $\pi^{E}$.
However, Example C.1 misleads the correct interpretation of the notion of reward function in MDPs, and in particular about the scale of the rewards. Let us explain better our point.
The MDP is a model, i.e., a simplified representation of reality, which is commonly applied to 2 different kinds of real-world scenarios: $(i)$ problems in which the agent (learner in RL or expert in IRL) actually receives some kind of scalar feedback from the environment, which can be modelled as a reward function; $(ii)$ problems in which the agent does not receive a feedback from the environment, but its objective, i.e., its structure of preferences among state-action trajectories (which trajectories are better than others), satisfies some axioms that permit to represent it through a scalar reward [52, 9] (this is referred to as the Reward Hypothesis in literature [58]).
There is an enormous difference between scenario $(i)$ and scenario $(ii)$. In $(i)$ the notion of $\epsilon$-optimal policy is well-defined for any fixed $\epsilon>0$, because the reward function is given and, thus, fixed. Instead, in $(ii)$, the notion of reward function is a mere mathematical artifact used to represent preferences among trajectories, whose existence is guaranteed by a set of assumptions/axioms [52, 9, 58]. As Example C.1 shows, positive affine transformations of the reward do not affect the structure of preferences represented (see [52] or Section 16.2 of [50] or [30]). Therefore, in $(ii)$, the notion of $\epsilon$-optimal policy is not well-defined, because rescaling a reward function $r$ to $kr$ changes the suboptimality of some policy $\pi$ from $\epsilon$ to $k\epsilon$. In other words, for fixed $\epsilon>0$, any policy can be made $\epsilon$-optimal by simply rescaling a reward $r$ to $kr$ for some small enough $k>0$.
In IRL, this issue is even more influential because, although we are in setting $(i)$, we have no idea on the scale of the true reward function. For this reason, our solution is to attach to any reward $r$ a notion of compatibility $\overline{\mathcal{C}}(r)$ which implicitly contains information about the scale of the reward $r$. Compatibilities of different rewards (e.g., $r^{1}$ and $r^{2}$ in Example C.1) cannot be compared unless the rewards have the same scale (e.g., $r^{1}$ and $r^{2}$ have different scales, thus their compatibilities shall not be compared).
It should be observed that in Appendix C.2 we discuss a notion of compatibility independent of the scale of the reward. However, we show that it suffers from major drawbacks that make the notion of compatibility introduced in the main paper (Definition 4.1) more suitable for the IRL problem.
In conclusion, to settle Example C.1, rewards $r^{1}$ and $r^{2}$ should not have the same compatibility, because they have different scales, and the notion of compatibility (i.e., suboptimality) is strictly connected to the scale of the reward. To carry out a fair comparison of compatibilities, one should rescale the compatibility of each reward based on the scale of the reward.